OpenStudy (anonymous):

Identify the 42nd term of an arithmetic sequence where a1 = –12 and a27 = 66. al=1st term, a27=27th term.

OpenStudy (anonymous):

General term of an arithmetic sequence for nth term is\[a_n = a_1+d(n-1)\]where d is the common difference. \[a_{27} = 66\]\[a_{27} =-12+d(27-1)=66\]\[∴d=3\]Now use the general term to find the 42nd term\[a_{42}=-12+3(42-1)=111\] You're welcome :)